If A + B + C = 180° then the value of tanA + tanB + tanC is : how to solve

    If  A + B + C = 180° then the value of tanA + tanB + tanC is : how to solve

Grade:12

1 Answers

Sudheesh Singanamalla
13 years ago

we have A+B=180-C


tan(A+B)=tan(180-C)
[tanA+tanB]/[1-tanAtanB] = -tanC
tanA + tanB = - tanC+ tanA * tanB * tanC
tanA + tanB + tanC = tanA * tanB * tanC

 

Please approve